# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v3

@@ VIII Conclusion
  • $\rho_{p_{i}}(t)=0.1+(0.5-0.1)e^{-0.5t}$, $i=\{x,y,z\}$,
  • $\rho_{v_{x}}(t)=0.5+(5-0.5)e^{-0.2t}$, $\rho_{v_{y}}(t)=0.5+(5-0.5)e^{-0.2t}$, $\rho_{v_{z}}(t)=0.2+(5-0.2)e^{-1.5t}$
  • $\rho_{\phi\theta_{1}}(t)=0.1+(0.5-0.1)e^{-0.5t}$, $\rho_{\phi\theta_{2}}(t)=0.1+(0.5-0.1)e^{-0.5t}$, $\rho_{\psi}(t)=0.1+(0.4-0.1)e^{-0.05t}$
  • $\rho_{w_{i}}(t)=0.3+(0.5-0.3)e^{-0.5t}$, $i=\{x,y,z\}$
  • $K_{p}=\mathrm{\textnormal{diag}}\!\left(1.25,1.25,12.5\right)$
  • $K_{v}=\mathrm{\textnormal{diag}}\!\left(20,20,5\right)$
  • $K_{\phi\theta}=\mathrm{\textnormal{diag}}\!\left(3,1.5\right)$, $K_{\psi}=1$,
  • $K_{w}=10\cdot\mathbf{I}_{3}$.
@@
However, despite extensive efforts, we could not find a universal hyper-parameter that effectively tracked KDF’s trajectories, which may be due to RRT-generated adversarial trajectories being difficult for the controller to follow. To address this, we used smoothing splines to fit the RRT plans to adhere to the twice-differentiability required by funnel controllers but still encountered challenges with universal hyper-parameter settings. Therefore, assuming the existence of an effective funnel controller for KDF, we evaluated its performance based on planned trajectories, and padding obstacles by $\rho_{p}(0)$ as per the MATLAB Navigation Toolbox’s basic RRT implementation. We also present a selection of successfully tracked trajectories by KDF in Fig. 11 to validate our implementation of KDF.
Figure 15: The plot of realized trajectory in narrow-gap scenario, clearance of $0.31m$ for PARC. The 306 reach-avoid plans were computed out of 3,750 initial states.
-G Additional Experiments
@@
-G1 Planning and Tracking Model
We consider the time-switched polynomial model defined by (85a), (85c), (86a), and (86c), with $t_{\mathrm{\textnormal{f}}}=2$, $t_{\mathrm{\textnormal{pk}}}=1.5$, the planning states and workspace $\mathbf{p}=\mathbf{w}=[p_{x},p_{z}]^{\intercal}$, and trajectory parameters $\mathbf{k}=[k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x},k_{vz},k_{az},k_{\mathrm{\textnormal{pk}}z}]^{\intercal}$. Just like in Section -E2, we separated the planning model into low-dimensional models for each of the two workspace dimensions, and recovered the high-dimensional reach set, avoid set, and BRAS by (87) and (88).
@@
-G3 Hypothesis
We anticipate PARC to be able to compute a less conservative, but still guaranteed safe BRAS compared to Section V-A due to the reduced number (three vs. zero) of non-ETI states in the planning model. Moreover, we also expect PARC’s computation time to be faster than the naïve model choice, as the dimension for polytopic computation has been reduced from 8 to 4.
Since the planning model in Appendix -G1 also incorporates stabilizing behavior for the quadrotor near the goal region [kousik2019safe, mueller2015computationally], we expect to produce similar reach-avoid results to [dawson2022safe], that is, the quadrotor would be able to safely navigate between the two obstacles while stopping at the goal region.
-G4 Results
@@
-H Drift Vehicle Demo Details
@@
$\mathbf{z}=\begin{bmatrix}p_{x}&p_{y}&\theta&\omega&v&\beta\end{bmatrix}^{\intercal}$ (95)
@@
$\begin{split}\dot{p}_{x}&=v\cos(\theta+\beta)\\ \dot{p}_{y}&=v\sin(\theta+\beta)\\ \dot{\theta}&=\omega\\ \dot{\omega}&=\frac{1}{I_{z}}(L_{a}{f_{y,\mathrm{\textnormal{f}}}}\cos(\delta)+L_{a}{f_{x,\mathrm{\textnormal{f}}}}\sin(\delta)-L_{b}{f_{y,\mathrm{\textnormal{r}}}})\\ \dot{v}&=\frac{1}{m}({f_{x,\mathrm{\textnormal{f}}}}\cos(\delta-\beta)-{f_{y,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)\\ &\qquad+{f_{x,\mathrm{\textnormal{r}}}}\cos(\beta)+{f_{y,\mathrm{\textnormal{r}}}}\sin(\beta))\\ \dot{\beta}&=\frac{1}{mv}({f_{x,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)-{f_{y,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)\\ &\qquad+{f_{x,\mathrm{\textnormal{r}}}}\cos(\beta)+{f_{y,\mathrm{\textnormal{r}}}}\sin(\beta))-\omega\end{split}$ (96)
@@
${f_{y,\mathrm{\textnormal{f}}}}$ $=f_{\text{front}}(\omega,v,\beta,{f_{x,\mathrm{\textnormal{f}}}},\delta)\quad\mathrm{\textnormal{and}}$ (97)
@@
are formulated using the Fiala brush tire model [fiala1954lateral]. Lastly, the car’s physical parameters ($I_{z},m,L_{a},L_{b}$) as well as other constants used in the Fiala tire model are adopted from the DeLorean car specs from [goh2016simultaneous].
-H2 Feedback Controller
@@
Now we design the parameters in (99). Unlike our other system models, due to the lack of controllability during drifting, the behavior of the realized drift parking trajectories are “open-loop” in that they are uniquely and entirely determined given $\mathbf{z}_{0}$ and $\mathbf{k}$ without needing to provide the planning model, as illustrated in Section -H2. As such, we fit a planning model to the data collected in open-loop simulations and use PARC to compute the BRAS of the initial states.
@@
We use the geometry specs of the 1981 DeLorean ($l=4.267,w=1.988$unitm) [goh2020toward, goh2016simultaneous] to represent the car agent as a 2D rectangular H-Polytope $\overline{B}$. The obstacles ${\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{1},{\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{2}$ are two parked car obstacles centered at $[24,-16,-\pi]^{\intercal}$ and $[37,-16,-\pi]^{\intercal}$ each with the same geometry as $\overline{B}$.
@@
$\mathbf{p}_{\mathrm{\textnormal{goal}}}=\begin{bmatrix}30.5&-16&-\pi\end{bmatrix}^{\intercal}$
@@
Using the drift parking test environment, we collect 10,000 expert trajectories generated by feeding samples of $\mathbf{k}\in K=[9,11]\times[0.6109,0.7854]$ into the drift parking controller described above. The initial condition of all trajectories $\mathbf{z}_{0}$ is all zeros except for the velocity state $v_{0}=0.01$, as the tracking model experiences numerical instability at low velocities and sideslip angles. The entire simulation process lasts for 2 hours. With the expert trajectories ready, we create the time-variant affine planning model using the formulation above.
Figure 17: We are preparing an F1:10 class robotic vehicle to implement and deploy our safe drifting maneuvers. We aim to have a working demo by May 2024.

### v4

@@ VIII Conclusion
$\rho_{p_{i}}(t)$ $=0.1+(0.5-0.1)e^{-0.5t},\quad i=\{x,y,z\},$ (90a)
$\rho_{v_{x}}(t)$ $=0.5+(5-0.5)e^{-0.2t},$ (90b)
$\rho_{v_{y}}(t)$ $=0.5+(5-0.5)e^{-0.2t},$ (90c)
$\rho_{v_{z}}(t)$ $=0.2+(5-0.2)e^{-1.5t},$ (90d)
$\rho_{\phi\theta_{1}}(t)$ $=0.1+(0.5-0.1)e^{-0.5t},$ (90e)
$\rho_{\phi\theta_{2}}(t)$ $=0.1+(0.5-0.1)e^{-0.5t},$ (90f)
$\rho_{\psi}(t)$ $=0.1+(0.4-0.1)e^{-0.05t},$ (90g)
$\rho_{w_{i}}(t)$ $=0.3+(0.5-0.3)e^{-0.5t},\quad i=\{x,y,z\},$ (90h)
$K_{p}$ $=\mathrm{\textnormal{diag}}\!\left(1.25,1.25,12.5\right),$ (90i)
$K_{v}$ $=\mathrm{\textnormal{diag}}\!\left(20,20,5\right),$ (90j)
$K_{\phi\theta}$ $=\mathrm{\textnormal{diag}}\!\left(3,1.5\right),$ (90k)
$K_{\psi}$ $=1,$ (90l)
$K_{w}$ $=10\cdot\mathbf{I}_{3}.$ (90m)
@@
However, despite extensive efforts, we could not find a universal hyper-parameter that effectively tracked KDF’s trajectories, which may be due to RRT-generated adversarial trajectories being difficult for the controller to follow. To address this, we used smoothing splines to fit the RRT plans to adhere to the twice-differentiability requirement of funnel controllers but still encountered challenges with universal hyper-parameter settings. Therefore, assuming the existence of an effective funnel controller for KDF, we evaluated its performance based on planned trajectories, and padding obstacles by $\rho_{p}(0)$ as per the MATLAB Navigation Toolbox’s basic RRT implementation. We also present a selection of successfully tracked trajectories by KDF in Fig. 11 to validate our implementation of KDF.
[…]
-H Additional Experiments
@@
-H1 Planning and Tracking Model
We consider the time-switched polynomial model defined by (85a), (85c), (86a), and (86c), with $t_{\mathrm{\textnormal{f}}}=2$, $t_{\mathrm{\textnormal{pk}}}=1.5$, the planning states and workspace $\mathbf{p}=\mathbf{w}=[p_{x},p_{z}]^{\intercal}$, and trajectory parameters $\mathbf{k}=[k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x},k_{vz},k_{az},k_{\mathrm{\textnormal{pk}}z}]^{\intercal}$. Just like in Section -F2, we separated the planning model into low-dimensional models for each of the two workspace dimensions, and recovered the high-dimensional reach set, avoid set, and BRAS by (87) and (88).
@@
-H3 Hypothesis
We anticipated PARC to be able to compute a less conservative, but still guaranteed safe BRAS compared to Section V-A due to the reduced number (three vs. zero) of non-ETI states in the planning model. Moreover, we also expected PARC’s computation time to be faster than the naïve model choice, as the dimension for polytopic computation has been reduced from 8 to 4.
Since the planning model in Appendix -H1 also incorporates stabilizing behavior for the quadrotor near the goal region [kousik2019safe, mueller2015computationally], we expected to produce similar reach-avoid results to [dawson2022safe], that is, the quadrotor would be able to safely navigate between the two obstacles while stopping at the goal region.
-H4 Results
@@
-I Drift Vehicle Demo Details
@@
$\mathbf{z}=[p_{x},p_{y},\theta,\omega,v,\beta]^{\intercal}$ (91)
@@
$\begin{split}\dot{p}_{x}=&v\cos(\theta+\beta)\\ \dot{p}_{y}=&v\sin(\theta+\beta)\\ \dot{\theta}=&\omega\\ \dot{\omega}=&\frac{1}{I_{z}}(L_{a}{f_{y,\mathrm{\textnormal{f}}}}\cos(\delta)+L_{a}{f_{x,\mathrm{\textnormal{f}}}}\sin(\delta)-L_{b}{f_{y,\mathrm{\textnormal{r}}}})\\ \dot{v}=&\frac{1}{m}({f_{x,\mathrm{\textnormal{f}}}}\cos(\delta-\beta)-{f_{y,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)\\ &+{f_{x,\mathrm{\textnormal{r}}}}\cos(\beta)+{f_{y,\mathrm{\textnormal{r}}}}\sin(\beta))\\ \dot{\beta}&=\frac{1}{mv}({f_{x,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)-{f_{y,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)\\ &+{f_{x,\mathrm{\textnormal{r}}}}\cos(\beta)+{f_{y,\mathrm{\textnormal{r}}}}\sin(\beta))-\omega\end{split}$ (92)
@@
${f_{y,\mathrm{\textnormal{f}}}}$ $=f_{\text{front}}(\omega,v,\beta,{f_{x,\mathrm{\textnormal{f}}}},\delta),\quad\mathrm{\textnormal{and}}$ (93)
@@
are formulated using the Fiala brush tire model [fiala1954lateral]. Lastly, the car’s physical parameters ($I_{z},m,L_{a},L_{b}$) as well as other constants used in the Fiala tire model are adopted from the DeLorean car specification from [goh2016simultaneous].
-I2 Feedback Controller
@@
Now we design the parameters in (95). Unlike our other system models, due to the lack of controllability during drifting, the behavior of the realized drift parking trajectories are “open-loop” in that they are uniquely and entirely determined given $\mathbf{z}_{0}$ and $\mathbf{k}$ without needing to provide the planning model, as illustrated in Section -I2. As such, we fit a planning model to the data collected in open-loop simulations and use PARC to compute the BRAS of the initial states.
@@
We use the geometry specification of the 1981 DeLorean ($l=4.267,w=1.988$unitm) [goh2020toward, goh2016simultaneous] to represent the car agent as a 2D rectangular H-Polytope $\overline{B}$. The obstacles ${\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{1},{\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{2}$ are two parked car obstacles centered at $[24,-16,-\pi]^{\intercal}$ and $[37,-16,-\pi]^{\intercal}$ each with the same geometry as $\overline{B}$.
@@
$\mathbf{p}_{\mathrm{\textnormal{goal}}}=[30.5,-16,-\pi]^{\intercal}$
@@
[…]
Figure 17: We are preparing an F1:10 class robotic vehicle to implement and deploy our safe drifting maneuvers.